# Gianluca Monaci          Rafael S. Rezende          Romain Deffayet          Gabriela Csurka Guillaume Bono          Hervé Déjean          Stéphane Clinchant          Christian Wolf

### v1

@@ Abstract
Methods for navigation based on large-scale learning typically treat each episode as a new problem, where the agent is spawned with a clean memory in an unknown environment. While these generalization capabilities to an unknown environment are extremely important, we claim that, in a realistic setting, an agent should have the capacity of exploiting information collected during earlier robot operations. We address this by introducing a new retrieval-augmented agent, trained with RL, capable of querying a database collected from previous episodes in the same environment and learning how to integrate this additional context information. We introduce a unique agent architecture for the general navigation task, evaluated on ObjectNav, ImageNav and Instance-ImageNav. Our retrieval and context encoding methods are data-driven and heavily employ vision foundation models (FM) for both semantic and geometric understanding. We propose new benchmarks for these settings and we show that retrieval allows zero-shot transfer across tasks and environments while significantly improving performance.
@@ 1 Introduction
A realistic setting for robot navigation is the “unboxing” scenario, where one robot (or a fleet) is placed in its target setting, started-up, and navigates “out of the box”. The main appeal of this scenario is the lack of dependency on environment preparation, like scanning the environment, installing external localization systems or generating a floor plan prior to navigation. Most modern methods tackling this scenario, e.g. [75, 12, 79, 6], use learning-based approaches in an episodic setting: every episode in the robot’s life is treated as if it was its first. This is obviously sub-optimal: during operation the scene is explored, objects and their locations are observed, affordances found, failure cases encountered etc. Exploiting this information is crucial for designing efficient navigation agents that can continuously improve.
Figure 1: We augment a navigation agent with context information retrieved from a scene-specific retrieval database, fed with data by one or potentially multiple robots. We leverage semantic and multi-modal foundation models for retrieval, and geometric foundation models for context encoding, harnessing its power to provide directional information. We explore ObjectNav, ImageNav and Instance-ImageNav tasks, including zero-shot cases.
@@
Benchmarks like the k-item scenario [4], multi-object navigation [73] or GOAT benchmark [39] test the capacity of an agent to retain information from previous (sub-)goals and trajectories. Structured latent memory trained end-to-end with RL performs best in these benchmarks [50]. However, these architectures have only been tested on short (multi-)episodes of $500$ to $2,500$ agent steps and are hardly suitable for a real continuous operation: transformer-based agents with self-attention over time suffer from limited context length and the quadratic complexity of attention, while recurrent models are limited by the size of their latent memory and the network capacity growing quadratically with representation size. Alternative ways to integrate information from initial rollouts are topological maps [63, 69], which require the estimation and generation of a structured graph model from a sequence of posed observations. The extension of these methods to multi-robot scenarios is anything but trivial.
@@
In this work we propose a simple data-driven approach to extend a state-of-the-art navigation agent [6] working out-of-the-box, with capabilities to store visual observations in a global indexed database, share them with a fleet, and retrieve and process them to improve performance during operation. The proposed architecture is general and can be used for different navigation tasks — we evaluate our work on ObjectNav, ImageNav and Instance-ImageNav. While the goal may be specified either as an image or a semantic category, depending on the task, the retrieved information is always visual first person views (FPVs). Computer vision is at the heart of the agent in multiple ways: data-driven retrieval mechanisms leverage vision foundation models (FM) like CLIP and DINOv2 to retrieve candidate views, while a geometric FM [74] integrates the retrieved context and provides directional information to the agent, cf. Fig. 1.
@@
This approach is scalable, as the dataset can potentially be very large and can be queried by any number of agents with sub-linear time algorithms. Retrieving from an image database is also considerably simpler and more flexible than using involved memory structures such as metric or topological maps. The latter require pose to register observations and synchronization to enable multi-agent data collection: unless pose estimates are absolute (which requires scene preparation), fusing new updates from multiple agents requires a jointly maintained map and distributed algorithms.
@@
We propose the following contributions: (i) new navigation tasks, where the environment is augmented with a retrieval database fed by one or multiple robots; (ii) a unified end-to-end architecture for Retrieval-Augmented Navigation, RANa, for multiple tasks; (iii) data-driven retrieval mechanisms that do not require any meta-data beyond raw FPVs; (iv) mechanisms that leverage retrieved information either for additional context, or to replace the goal on the fly, allowing for zero-shot applications, cf. Fig. 2.
Figure 2: We target several contributions: perf. improvement (Success +x), opening new use cases from existing models, and zero-shot applications with retrieval only added at test time (Refer to caption).
Figure 3: We propose an architecture for Retrieval-Augmented Navigation architecture – RANa, that addresses all use cases of Fig. 2. Dashed arrows are optional depending on the navigation task.
@@ 2 Related Work
Visual navigation – The task of navigation has been addressed in robotics research using mapping and planning [71, 46], which requires solutions for mapping and localization [9, 44, 54], planning [41, 66] and low-level control [26, 62]. These methods depend on accurate sensor models, filtering, dynamical models and optimization schemes. In contrast, end-to-end models learn deep representations such as flat recurrent states [79, 6], occupancy maps [13], semantic maps [12], latent metric maps [56, 4, 32, 50], topological maps [3, 14, 67], self-attention [16, 23, 25, 61] or implicit representations [51]. Then, they map these representations into actions, using RL [36, 53, 8], imitation learning [21] or by maximizing navigability [7].
In this work, our goal is to address end-to-end learning of representations and policies with RL. Although this problem is nearly solved for PointNav, where the agent must navigate to a relative GPS coordinate [64, 75], it remains a challenge for other goal specifications. ObjectNav, where the goal is an object category, requires semantic understanding of the scene, therefore a large body of research builds a semantic map [27, 12] or leverages an LLM to guide navigation based on a scene’s common sense semantics [83, 10, 43, 82]. Several approaches address this task in a zero-shot fashion, training their agents without ObjectNav rewards and leveraging CLIP encoders for goal representation [70, 47, 27]. Recently, Zeng et al. [84] tackles ObjectNav with a transformer state encoder on simulated environments.
@@
In ImageNav, an image is provided as goal, requiring a geometric understanding of the scene. A common variant is Instance-ImageNav [42], where the goal image depicts a specific object in the environment and is taken with camera parameters different from those of the agent. Recently, the end-to-end model DEBiT [6] achieved state-of-the-art performance in ImageNav leveraging the geometric FM CroCo [74] to estimate relative poses between goal and agent’s observation. In this work, we build upon DEBiT by both extending it to ObjectNav and augmenting it with retrieval, naturally extending it to the multi-robot setting and enabling zero-shot ObjectNav and Instance-ImageNav.
To the best of our knowledge no other method performs all these tasks competently, the closest competitor being [47] which tackles ImageNav and zero-shot ObjectNav, although with considerably lower performance (see Sec. 5).
Method and Img Obj Not req.: Easiness of Multiple
representation Nav Nav pose depth repr. update collectors
GOAT [39] $\sim$ map reg.
3D Scene graphs [82] $\sim$ graph+map reg.
Topo map [63, 14, 67] $\sim$ graph upd.
Semantic maps [12] $\sim$ map reg.
Sem. Implicit [51] $\nabla$ upd
Feature SLAM [54] $\sim$ pose graph
Dense SLAM [24] $\sim$ map reg.
@@
To support multiple collectors, a method is either pose free, like RANa, or uses distributed algorithms to synchronize, or absolute pose (and not episodic).
@@
Retrieval-augmented control – The idea of retrieving information from a dataset of past experiences has been introduced in earlier works. Among them, episodic control methods for RL [5, 57, 45, 29, 30, 33] act on successful experiences by re-employing Q-value estimates, enhancing sample efficiency during training. In contrast, we exploit retrieved information from past experiences of a robot fleet to enhance navigation performance and to enable zero-shot task generalization. Goyal et al. [28] augment an RL agent with a retrieval process parameterized as a neural network that has access to a dataset of past trajectories. Here instead we take inspiration from [34] that, to our knowledge, was the first work to propose a generic non-parameterized retrieval augmentation for an RL agent with previously learned embeddings. However, alleviating the need to pre-train embeddings using RL, we use vision foundation models to process stored and retrieved images. This enables zero-shot transfer and allows to adapt the approach to various navigation tasks.
Retrieval-Augmented Planning [38] uses a memory of successful task executions, comprising plans, actions and observation sequences. Past experiences relevant to the task at hand are leveraged to guide planning with an LLM agent based on task constraints and state similarity. Xie et al. [78] build a hierarchical topological memory to prompt an LLM for reasoning and goal-finding in a navigable environment. In contrast, we only store images derived from unstructured observations without any associated metadata.
Navigation supported by prior rollouts – Other navigation methods exploit previous observations to construct a topological map of the explored scene, and use it to guide navigation [63, 14, 67]. In contrast, our database is unstructured and we can query it to retrieve goal and context images with simple image retrieval.
In Tab. 1 we compare RANa with various navigation methods discussed above in the continual navigation scenario. Most approaches require depth and pose and/or more involved (metric or topological) map updates, which need at least to register observations. In contrast, RANa does not require depth or pose, the database update and retrieval is very simple, and naturally supports multiple data collection robots without the need of synchronization. While SLAM-based solutions [24, 54] are well established for continual navigation, with RANa we propose a lightweight data-driven solution that can be easily scaled up to take inputs from multiple robots and can tackle ImageNav, Instance-ImageNav and ObjectNav use cases.
@@ 3 Retrieval-augmented agent
We target navigation in 3D environments, where an agent is tasked to navigate from a starting location to a goal and receives at each timestep $t$ an image observation $\mathbf{x}_{t}$. Our method is general and the agent can be trained for a diverse range of tasks, but without loss of generality in this work we focus on ImageNav, where the agent receives an additional image goal and ObjectNav, where the agent receives an object goal which is a category in form of text, cf. Fig. 2. In both cases we denote the original goal by $\mathbf{g}$.
The action space is discrete, $\mathcal{A}$ ={MOVE FORWARD 0.25m, TURN LEFT $10^{\circ}$, TURN RIGHT $10^{\circ}$, and STOP}. Navigation is considered successful if the STOP action is selected when the agent is within 1m of the goal position in terms of geodesic distance. ObjectNav has two additional actions, LOOK UP and LOOK DOWN.
Our tasks are based on the ImageNav and ObjectNav tasks of the Habitat simulator and platform [64]. We extended them with a retrieval dataset $\mathcal{D}$ and a retrieval mechanism $\mathcal{R}$, which will be detailed in Sec. 4. The dataset $\mathcal{D}=\{\mathbf{x}_{i}^{\mathcal{D}}\}$ contains FPVs indexed by $i$ and stemming from previous episodes or from exploratory rollouts in the scene. While it is in principle possible to store additional meta-data associated to data collection, like approximate pose, performed actions, reward, success or value (in an MDP sense, when available), we show in this work that the access to FPVs alone provides rich information to the agent (see Sec. 4).
The agent queries this dataset $\mathcal{D}$ with the retrieval mechanism $\mathcal{R}$, which can be potentially used at each time step $t$. $\mathcal{R}$ uses the goal $\mathbf{g}$, and optionally the current observation $\mathbf{x}_{t}$, to retrieve a set of FPVs $\mathcal{R}(\mathbf{x}_{t},\mathbf{g})\longrightarrow\{\mathbf{r}_{t,1},\mathbf{r}_{t,2},\dots\mathbf{r}_{t,N}\}$, that is used as context information to improve navigation. Depending on the use case, the use of observation $\mathbf{x}_{t}$ in the retrieval process may be optional — if not provided, retrieval can be done once per episode, as opposed to every time step.
@@
$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// compare obs+goal}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// update rec state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// act}}\\ \end{array}$
where $x$ and $g$ are trainable encoders, $h$ is the update function of a GRU [18] which maintains a hidden state $\mathbf{h}_{t}$ over time; $l$ is an embedding function, and $\pi$ is the policy. For clarity we have omitted the equations of the gating functions.
@@
$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// compare obs+goal}}\\ \tilde{\mathbf{c}}_{t}&=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// retrieve context}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},\tilde{\mathbf{c}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// update rec state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// act}}\\ \end{array}$
@@
where $c$ is an encoder which compresses the retrieved context into a compact form. It is trained end-to-end together with the agent, as described in Sec. 3.1, while the retrieval process $\mathcal{R}$, described in Sec. 4, leverages pre-trained models. In all tasks, the observation encoder $x$ is implemented as a half-width ResNet-18 (hwRN18) architecture [31].
The goal $\mathbf{g}$, which is an image or a category, according to the navigation task, is compared to the observation $\mathbf{x}_{t}$ through a dedicated comparison function $g$ defined as
$g(\mathbf{x}_{t},\mathbf{g})=\left\{\begin{array}[]{ll}d(\mathbf{x}_{t},\mathbf{g})&{\scriptsize\texttt{if ImageNav}}\\ [cl_{V}(\mathbf{x}_{t}),cl_{T}(\mathbf{g})]&{\scriptsize\texttt{if ObjectNav}}\\ \end{array}\right.$ (3)
@@
As in [6], for ImageNav we use a binocular visual encoder $d$ that compares the observation $\mathbf{x}_{t}$ with the goal image $\mathbf{g}$ and returns an embedding representing information about the goal’s direction, pose, and visibility. This encoder $d$ has been fine-tuned from the foundation model CroCo [74], solving as pre-text tasks relative-pose and visibility estimation [6]. We use the weights available online. For ObjectNav where a goal category is specified, we use a CLIP encoder [59] providing a semantic comparison of the observed image, encoded as $cl_{V}(\mathbf{x}_{t})$, and the object category, encoded as $cl_{T}(\mathbf{g})$.
@@
Additionally, an image goal $\mathbf{g}^{\mathcal{R}}$ can be retrieved from $\mathcal{D}$ given the task goal $\mathbf{g}$, allowing the ImageNav model to address tasks in a zero-shot way. For example, an ImageNav agent can perform ObjectNav by retrieving a goal image $\mathbf{g}^{\mathcal{R}}$ from the database $\mathcal{D}$ given a goal category ([Uncaptioned image], [Uncaptioned image]). This can be integrated into the augmented agent above by adding a third case to the goal comparison function $g$,
$g(\mathbf{x}_{t},\mathbf{g})=d(\mathbf{x}_{t},\mathbf{g}^{\mathcal{R}})\leavevmode\nobreak\ \leavevmode\nobreak\ {\scriptsize\texttt{if image goal isretrieved}}$ (4)
We explore this use of retrieval for zero-shot task generalization in Sec. 5.2.

### v2

# RANa: Retrieval-Augmented Navigation

@@ Abstract
Methods for navigation based on large-scale learning typically treat each episode as a new problem, where the agent is spawned with a clean memory in an unknown environment. While these generalization capabilities to an unknown environment are extremely important, we claim that, in a realistic setting, an agent should have the capacity of exploiting information collected during earlier robot operations. We address this by introducing a new retrieval-augmented agent, trained with RL, capable of querying a database collected from previous episodes in the same environment and learning how to integrate this additional context information. We introduce a unique agent architecture for the general navigation task, evaluated on ImageNav, Instance-ImageNav and ObjectNav. Our retrieval and context encoding methods are data-driven and employ vision foundation models (FM) for both semantic and geometric understanding. We propose new benchmarks for these settings and we show that retrieval allows zero-shot transfer across tasks and environments while significantly improving performance.
Refer to caption(a) agent(b) use cases(c) inference runtime
Figure 1: Retrieval-Augmented Navigation (RANa). (a) We augment a navigation agent with context information retrieved from a scene-specific retrieval database, fed with data by one or potentially multiple robots. We leverage semantic or multi-modal foundation models for retrieval, and geometric foundation models for context encoding. (b) We tackle ImageNav, Instance-ImageNav and ObjectNav tasks, demonstrating performance improvements (Success +x), opening new use cases from existing models, and zero-shot applications with retrieval only added at test time (Refer to caption). (c) Retrieval is efficient, neglectable compared to vision models and keeps inference run-time in the same ballpark as non-augmented models like DEBiT.
@@ 1 Introduction
A realistic setting for robot navigation is the “unboxing” scenario, where one robot (or a fleet) is placed in its target setting, started-up, and navigates “out of the box”. The main appeal of this scenario is the lack of dependency on environment preparation, like scanning the environment, installing external localization systems or generating a floor plan prior to navigation. Most modern methods tackling this scenario, e.g. Wijmans et al. (2019); Chaplot et al. (2020a); Yadav et al. (2023a); Bono et al. (2024a), use learning-based approaches in an episodic setting: every episode in the robot’s life is treated as if it was its first. This is obviously sub-optimal: during operation the scene is explored, objects and their locations are observed, affordances found, failure cases encountered etc. Exploiting this information is crucial for designing efficient navigation agents that can continuously improve.
@@
Benchmarks like the k-item scenario (Beeching et al., 2020b), multi-object navigation (Wani et al., 2020) or GOAT benchmark (Khanna et al., 2024) test the capacity of an agent to retain information from previous (sub-)goals and trajectories. Structured latent memory trained end-to-end with RL performs best in these benchmarks (Marza et al., 2022). However, these architectures have only been tested on short (multi-)episodes of 500500500 to 2,5002,5002 , 500 agent steps and are hardly suitable for a real continuous operation: transformer-based agents with self-attention over time suffer from limited context length and the quadratic complexity of attention, while recurrent models are limited by the size of their latent memory and the network capacity growing quadratically with representation size (Jose et al., 2018). Alternative ways to integrate information from initial rollouts are topological maps (Savinov et al., 2018; Sridhar et al., 2024), which require the estimation and generation of a structured graph model from a sequence of posed observations. The extension of these methods to multi-robot scenarios is anything but trivial.
@@
In this work we propose a simple data-driven approach to extend a state-of-the-art navigation agent (Bono et al., 2024a) working out-of-the-box, with capabilities to store visual observations in a global indexed database, share them with a fleet, and retrieve and process them to improve performance during operation. The proposed architecture is general and can be used for different navigation tasks — we evaluate our work on ImageNav, Instance-ImageNav and ObjectNav. While the goal may be specified either as an image or a semantic category, depending on the task, the retrieved information is always visual first person views (FPVs). Data-driven retrieval mechanisms leverage vision foundation models (FM) like DINOv2 (Oquab et al., 2024) to retrieve candidate views, while a geometric FM, CroCo (Weinzaepfel et al., 2022), integrates the retrieved context and provides directional information to the agent, cf. Fig.˜1a. Retrieval is also fast and keeps inference time in the same ball park of alternative baselines, cf. Fig.˜1c.
@@
This approach is scalable, as the dataset can potentially be very large and can be queried by any number of agents with optimized sub-linear time algorithms. Retrieving from an image database is also considerably simpler and more flexible than using involved memory structures such as metric or topological maps. While we can use any suitable vision or multi-modal FM to represent, organize and retrieve images from an unstructured database, the latter approaches require pose to register observations and synchronization to enable multi-agent data collection: unless pose estimates are absolute (which requires scene preparation), fusing new updates from multiple agents requires a jointly maintained map and distributed algorithms.
@@
We propose the following contributions: (i) new navigation tasks, where the environment is augmented with a retrieval database fed by one or multiple robots; (ii) a unified end-to-end architecture for Retrieval-Augmented Navigation, RANa, for multiple tasks; (iii) data-driven retrieval mechanisms that do not require any meta-data beyond raw FPVs; (iv) mechanisms that leverage retrieved information either for additional context, or to replace the goal on the fly, allowing for zero-shot applications, cf. Fig.˜1.
@@ 2 Related Work
Visual navigation – The task of navigation has been addressed in robotics research using mapping and planning (Thrun et al., 2005; Macenski et al., 2020), which requires solutions for mapping and localization (Bresson et al., 2017; Mur-Artal & Tardós, 2017; Labbé & Michaud, 2019), planning (Sethian, 1996; Konolige, 2000) and low-level control (Fox et al., 1997; Rösmann et al., 2015). These methods depend on accurate sensor models, filtering, dynamical models and optimization schemes. In contrast to them, end-to-end models aim to learn deep representations such as flat recurrent states (Yadav et al., 2023a; Bono et al., 2024a), occupancy maps (Chaplot et al., 2020b), semantic maps (Chaplot et al., 2020a), latent metric maps (Henriques & Vedaldi, 2018; Parisotto & Salakhutdinov, 2018; Beeching et al., 2020b; Marza et al., 2022), topological maps (Savinov et al., 2018; Beeching et al., 2020a; Chaplot et al., 2020c; Sridhar et al., 2024), self-attention (Fang et al., 2019; Du et al., 2021; Chen et al., 2022; Reed et al., 2022; Zeng et al., 2025) or implicit representations (Marza et al., 2023; Kwon et al., 2023). Then, they map these representations into actions, using RL (Jaderberg et al., 2017; Mirowski et al., 2017; Bono et al., 2024c), imitation learning (Ding et al., 2019) or by maximizing navigability (Bono et al., 2024b).
In this work, our goal is to address end-to-end learning of representations and policies with RL. Although this problem is nearly solved for PointNav, where the agent must navigate to a relative GPS coordinate (Savva et al., 2019; Wijmans et al., 2019), it remains a challenge for other goal specifications.
@@
Here we focus on ImageNav, where an image is provided as goal, requiring geometric understanding of the scene. A common variant is Instance-ImageNav (Krantz et al., 2023), where the goal image depicts a specific object in the environment and is taken with camera parameters different from those of the agent. Recently, the end-to-end model DEBiT (Bono et al., 2024a) achieved state-of-the-art performance in ImageNav leveraging the geometric FM CroCo (Weinzaepfel et al., 2022) to estimate relative poses between goal and agent’s observation. In this work, we build upon DEBiT and augment it with retrieval, naturally extending it to multi-robot setting and enabling zero-shot ObjectNav and Instance-ImageNav. To the best of our knowledge no other method performs all these tasks competently, the closest competitor being Majumdar et al. (2022) which tackles ImageNav and zero-shot ObjectNav, although with considerably lower performance (see Section˜5).
Retrieval-augmented control – The idea of retrieving information from a dataset of past experiences has been introduced in earlier works. Among them, episodic control methods for RL (Blundell et al., 2016; Pritzel et al., 2017; Lin et al., 2018; Hansen et al., 2019; He et al., 2024; Hu et al., 2021) act on successful experiences by re-employing Q-value estimates, enhancing sample efficiency during training. In contrast, we exploit retrieved information from past experiences of a robot fleet to enhance navigation performance and enable zero-shot task generalization. Goyal et al. (2022) augment an RL agent with a retrieval process parameterized as a neural network that has access to a dataset of past trajectories. Here instead, we take inspiration from Humphreys et al. (2022) that, to our knowledge, was the first work to propose a generic non-parameterized retrieval augmentation for an RL agent with previously learned embeddings. Alleviating the need to pre-train embeddings with RL, we use vision foundation models to process stored and retrieved images. This enables zero-shot transfer and allows to adapt the approach to various navigation tasks.
Retrieval-Augmented Planning (Kagaya et al., 2024) uses a memory of successful task executions, comprising plans, actions and observation sequences. Past experiences relevant to the task at hand are leveraged to guide planning with an LLM agent based on task constraints and state similarity. Xie et al. (2024) build a hierarchical topological memory to prompt an LLM for reasoning and goal-finding in a navigable environment. In contrast, we only store images derived from unstructured observations without any associated metadata.
Navigation supported by prior rollouts – Other navigation methods exploit previous observations to construct a topological map of the explored scene, and use it to guide navigation (Savinov et al., 2018; Chaplot et al., 2020c; Shah & Levine, 2022; Sridhar et al., 2024). In contrast, our database is unstructured and we can query it to retrieve goal and context images with simple image retrieval. Also, here context images are not used as navigation waypoints, but as recommendations, exploited or not by the agent, leading to navigation performance robust to potentially misleading context, as shown in Section˜5.
In Table˜1 we compare RANa with selected navigation methods discussed above in the continual navigation scenario. Most approaches require depth and pose and/or more involved (metric or topological) map updates, which need at least to register observations. In contrast, RANa does not require depth or pose, the database update and retrieval is straightforward, and naturally supports multiple data collection robots without the need of synchronization. While SLAM-based solutions (Engel et al., 2014; Mur-Artal & Tardós, 2017) are well established for continual navigation, with RANa we propose a lightweight data-driven solution that can be easily scaled up to multiple robots and can tackle ImageNav, Instance-ImageNav and ObjectNav use cases.
Method and Img Obj Not required: Easiness of Multiple
representation Nav Nav pose depth representation update collectors
GOAT (Khanna et al., 2024) $\sim$ map reg.
3D Scene graphs (Yin et al., 2024) $\sim$ graph+map reg.
Topo map (Savinov et al., 2018; Shah & Levine, 2022) $\sim$ graph update
Semantic maps (Chaplot et al., 2020a) $\sim$ map reg.
Sem. Implicit (Marza et al., 2023) $\nabla$ update
Feature SLAM (Mur-Artal & Tardós, 2017) $\sim$ pose graph
Dense SLAM (Engel et al., 2014) $\sim$ map reg.
@@
In order to support multiple collectors, a method should either be pose free, like RANa, or use distributed algorithms to synchronize, or absolute pose (and not episodic).
@@ 3 Retrieval-augmented agent
We target navigation in 3D environments, where an agent is tasked to navigate from a starting location to a goal and receives at each timestep $t$ an image observation $\mathbf{x}_{t}$. Our method is general and the agent can be trained for a diverse range of tasks, but without loss of generality in this presentation we focus on ImageNav, where the agent receives a goal image $\mathbf{g}$. An extension to the ObjectNav task is presented in Section˜6. The action space is discrete, $\mathcal{A}$ ={MOVE FORWARD 0.25m, TURN LEFT $10^{\circ}$, TURN RIGHT $10^{\circ}$, and STOP}. Navigation is considered successful if the STOP action is selected when the agent is within 1m of the goal position in terms of geodesic distance.
Our set-up is based on the ImageNav task of the Habitat simulator and platform (Savva et al., 2019). We extend it with a retrieval dataset $\mathcal{D}$ and a retrieval mechanism $\mathcal{R}$, which will be detailed in Section˜4. The dataset $\mathcal{D}=\{\mathbf{x}_{i}^{\mathcal{D}}\}$ contains FPVs indexed by $i$ and stemming from previous episodes or from exploratory rollouts in the scene. While it is in principle possible to store additional meta-data associated to data collection, like approximate pose, performed actions, reward, success or value (in an MDP sense, when available), in Section˜5 we show that access to FPVs alone provides rich, useful information to the agent.
The agent queries this dataset $\mathcal{D}$ with the retrieval mechanism $\mathcal{R}$, which can be potentially used at each time step $t$. $\mathcal{R}$ uses the goal $\mathbf{g}$, and optionally the current observation $\mathbf{x}_{t}$, to retrieve a set of FPVs $\mathcal{R}(\mathbf{x}_{t},\mathbf{g})\longrightarrow\{\mathbf{r}_{t,1},\mathbf{r}_{t,2},\dots\mathbf{r}_{t,N}\}$, that is used as context information to improve navigation. Depending on the use case, the use of observation $\mathbf{x}_{t}$ in the retrieval process may be optional (as in the static context in Section˜4) — if not provided, retrieval can be done once per episode, as opposed to every time step.
@@
$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// compare obs+goal\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// update recurrent state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// act}}\\ \end{array}$
where $x$ and $g$ are trainable encoders, $h$ is the update function of a GRU (Cho et al., 2014) which maintains a hidden state $\mathbf{h}_{t}$ over time; $l$ is an embedding function, and $\pi$ is the policy. For clarity we have omitted the equations of the gating functions.
@@
$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// compare obs+goal}}\\ \tilde{\mathbf{c}}_{t}&=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// retrieve and encode context}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},\tilde{\mathbf{c}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// update recurrent state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// act}}\\ \end{array}$
@@
where $c$ is an encoder which compresses the retrieved context into a compact form. It is trained end-to-end together with the agent, as described in Section˜3.1, while the retrieval process $\mathcal{R}$, described in Section˜4, leverages pre-trained models. The observation encoder $x$ is implemented as a half-width ResNet-18 (hwRN18) architecture (He et al., 2016).
@@
The image goal $\mathbf{g}$ is compared to the observation $\mathbf{x}_{t}$ through a function $g(\mathbf{x}_{t},\mathbf{g})$ implemented with a binocular visual encoder from Bono et al. (2024a) that returns an embedding representing information about the goal’s direction, pose, and visibility. This encoder was fine-tuned from the foundation model CroCo (Weinzaepfel et al., 2022), solving as pre-text tasks relative-pose and visibility estimation.
@@
Additionally, an image goal $\mathbf{g}^{\mathcal{R}}$ can be retrieved from $\mathcal{D}$ given the task goal $\mathbf{g}$, allowing the ImageNav model to address tasks in a zero-shot way. For example, an ImageNav agent can perform ObjectNav by retrieving a goal image $\mathbf{g}^{\mathcal{R}}$ from the database $\mathcal{D}$ given a goal category ([Uncaptioned image], [Uncaptioned image]). In this case, the goal comparison function $g$ is computed between $\mathbf{x}_{t}$ and $\mathbf{g}^{\mathcal{R}}$ (instead of the original goal $\mathbf{g}$) as $g(\mathbf{x}_{t},\mathbf{g}^{\mathcal{R}})$. We explore this use of retrieval for zero-shot task generalization in Section˜5.2.
Figure 2: We propose an architecture for Retrieval-Augmented Navigation architecture – RANa, that addresses all use cases of Fig.˜1b. Dashed arrows are optional depending on the navigation task.